OpenStudy (anonymous):

Solve the following equation by transforming it into a perfect square trinomial: x^2 – 14x = 15

OpenStudy (primeralph):

you know the drill. Make an attempt.

OpenStudy (anonymous):

49?

OpenStudy (primeralph):

@Nurali you know the rules. Stop posting answers.

OpenStudy (anonymous):

I thought it was ok just as long as you show how to do it

OpenStudy (primeralph):

No, we're not allowed to show answers unless the asker makes some true effort.

OpenStudy (anonymous):

oh

OpenStudy (nurali):

x² - 14x = 15 x² - 7x = 15 + (- 7)² x² - 7x = 15 + 49 (x - 7)² = 64 x - 7 = 8 x = 15 x = - 8 + 7, x = - 1 Answer: - 1, 15

OpenStudy (primeralph):

@Nurali you don't listen do you?